OpenStudy (anonymous):

What are the vertex and the axis of symmetry of the equation? y = –2x2 + 8x – 18 (1 point) vertex: ( 2, –10) axis of symmetry: x = 2 vertex: (2, –10) axis of symmetry: x = –10 vertex: (–2, –10) axis of symmetry: x = –2 vertex: (–2, 10) axis of symmetry: y = –2

OpenStudy (anonymous):

h=-b/2a x=h is axis of symmetry (h,k) is vertex

OpenStudy (anonymous):

i picked c vertex: (–2, –10) axis of symmetry: x = –2 and it waz wrong yo

OpenStudy (anonymous):

a=-2 b=8 -b/2a = -8/-4 = 2.

OpenStudy (anonymous):

Check doz negative signs, yo.

OpenStudy (anonymous):

cool thanks